OpenStudy (anonymous):

Consider a solid orientated along the x axis such that its cross-sectional area is given by A(x) = e^(5 x) Find the volume of the solid contained between x = 17 and x = 23.

OpenStudy (dumbcow):

\[\large V = \int\limits_{17}^{23}A(x) dx\] \[\large = \int\limits_{17}^{23} e^{5x} dx\]

OpenStudy (anonymous):

\[\int\limits_{17}^{23}e ^{5x}\]

OpenStudy (dumbcow):

use substitution u = 5x du = 5 dx --> du/5 = dx \[\large \frac{1}{5}\int\limits_{17}^{23}e^{u} du = \frac{1}{5}(e^{5*23} - e^{5*17})\]